APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday April 29, 2003 (MJD 52759) ---=== OBSERVING TEAM ===--- Swing: Pete Newman Night: Steph Snedden, Mike Harvanek Visitors: John Stocke, Brian Keeney (University of Colorado) ---=== OBSERVING PLAN ===--- Science. ---=== OBSERVING SUMMARY ===--- Five plates finished (1016, 1282, 1320, 1336, 1172) and good data obtained for another plate (1239). ---=== OBSERVING LOG ===--- Swing shift notes ----------------- The HgCd lamp above counterweight 2 coninues to be slow to turn on. That above CW4 (SW corner) seemed a little slow, too. A 71-frame bias drift in the (lit) enclosure archived without incident. Night ----- Partly cloudy and borderline high dust counts at sunset. We left the doors closed and fans off while waiting for the dust to settle. We did not switch instruments while waiting for the dust to drop in hopes that the clouds would clear after sunset as they did last night. They did not and we so we switched to spectroscopy. Plate 1239 (cartridge 3): Needed an FK5 star to acquire guide stars but this was probably due to clouds rather than pointing. The FK5 star came and went with the clouds but the guide stars were finally acquired. Problems with the clouds continued, however. An exposure was attempted but was seriously affected by clouds. When this exposure ended, it never readout. An endStare was issued but did not work. An endStare -abort appeared to work at first but when we tried a bias to test things, we got messages about sp1 not being idle. Steph then found that although she thought she had clicked on the grabInst button in the sop gui, the command had never been issued and so the instruments were never grabbed. This is what caused the exposure to not read out. The exposure was finally read out using an endStare -force. At this point we realized we had forgotten to do a preCal (it must have been all that imaging the past 5 nights). However, looking at the exposure on the monitors, the data were not good anyway. The clouds started to thin at about this time and so we started over (preCal, etc.). We obtained one 15 minute exposure on this plate before it expired. A note regarding spectro exposure numbers is in order here. Son-of-spectro reports a mismatch of exposure numbers for the exposure that would not read out: WARNING: Exposure number in header (20644) disagrees w/filename (20640) and exposures 20640 through 20644 are missing from S-o-s. We believe the 5 missing exposures can be accounted for as follows: the original readout, the endStare, the endStare -abort, the test bias and the endStare -force. Since S-o-s did not complain about the next exposure (preCal flat 20645) we believe everything is ok. Plate 1016 (cartridge 5): After offsets were cleared, we still needed to hunt for the glare from an FK5 star. Plate finished in 3 exposures (2100 secs total). There was a weird ``spike'' in the cloud camera around 0700. It is uncertain whether or not this was a real sky event. Maybe there is an electronics problem or a problem with the chopper wheel. Perhaps someone should have a look? Although clear overhead, there were still clouds not far to the west and so we continued with spectroscopy. Plate 1282 (cartridge 6): Guide stars on the gotoField. Plate finished in 3 exposures (1800 secs total). Although the sky is clear, there are fast moving clouds from the west on the satellite images and there is some doubt that it will remain photometric long enough to make a switch to imaging worth while. So, we continue with spectroscopy. Plate 1320 (cartridge 4): Guide stars on the gotoField. Plate finished in 3 exposures (1800 secs total). Plate 1336 (cartridge 1): Guide stars on the gotoField. Plate finished in 3 exposures (1800 secs total). Plate 1172 (cartridge 7): Guide stars on the gotoField. When combined with previous data (52756), this plate was finished in 2 exposures (1320 secs total). endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 3921 20:56Z 21:35Z 100 O -119.34 -109.62 79 ignore 3922 01:16Z 01:33Z 100 O -54.03 -49.88 41 bias on telescope ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 20637, plate -9999 ------- 20:58 20637 0.0 bias 02:53 20639 0.0 bias Missing exposure numbers here. See main body of log for explanation. ----- sequence 20644, plate 1239 ------- 6.7 6.6 6.5 7.2 05:04 20644 1500.1 ignore 05:11 20645 10.0 flat 05:14 20646 2.0 arc 05:33 20647 900.1 target ----- sequence 20648, plate 1016 ------- 19.1 18.0 15.2 16.8 DONE 05:54 20648 10.0 flat 05:57 20649 2.0 arc 06:16 20650 900.1 target 06:30 20651 600.1 target 06:43 20652 600.1 target 06:54 20653 240.0 smear ----- sequence 20654, plate 1282 ------- 23.1 21.2 20.8 23.4 DONE 07:07 20654 10.0 flat 07:09 20655 2.0 arc 07:24 20656 600.1 target 07:37 20657 600.1 target 07:49 20658 600.1 target 07:57 20659 240.0 smear ----- sequence 20660, plate 1320 ------- 16.4 16.1 14.6 17.8 DONE 08:16 20660 10.0 flat 08:18 20661 2.0 arc 08:33 20662 600.1 target 08:46 20663 600.1 target 08:59 20664 600.1 target 09:08 20665 240.0 smear ----- sequence 20666, plate 1336 ------- 20.3 15.9 19.0 19.4 DONE 09:19 20666 10.0 flat 09:22 20667 2.0 arc 09:37 20668 600.1 target 09:50 20669 600.1 target 10:03 20670 600.1 target 10:11 20671 240.0 smear ----- sequence 20672, plate 1172 ------- 4.9 3.8 4.3 3.3 52756 10:23 20672 10.0 flat 14.6 14.8 14.5 15.4 tonight 10:26 20673 2.0 arc ------------------- 10:42 20674 720.1 target 19.5 18.6 18.8 18.7 DONE 10:50 20675 240.0 smear 11:03 20676 600.1 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 04:20Z 3 1239 -41.70 0.0045 59.96 0.0008 145.45 0.0000 1.000240 06:05Z 5 1016 212.01 0.0031 61.33 0.0026 50.54 -0.0204 1.000190 07:22Z 6 1282 218.73 0.0031 67.04 0.0026 53.17 -0.0204 1.000170 08:26Z 4 1320 215.95 0.0009 55.10 0.0010 63.20 -0.0036 1.000030 09:28Z 1 1336 230.28 0.0010 79.03 0.0014 56.61 -0.0056 1.000080 10:29Z 7 1172 242.66 0.0025 70.18 0.0009 76.36 -0.0109 1.000150 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL6367 Stays: JL6374 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 04:19Z 3 1239 1.00024 -2369 -1035 -225 -41 60.1 9.9 20 268 ---- 2.0" 06:05Z 5 1016 1.00019 -1875 -579 -175 212 61.4 8.7 14 275 ---- 1.7" 07:21Z 6 1282 1.00017 -1678 -378 -125 219 67.1 7.6 25 268 ---- 1.6" 08:26Z 4 1320 1.00003 -296 870 -50 216 55.2 7.0 22 270 ---- 1.5" 09:29Z 1 1336 1.00008 -789 390 -75 230 79.0 6.6 13 272 ---- 1.5" 10:29Z 7 1172 1.00015 -1480 -195 -100 243 70.1 5.7 26 269 ---- 1.4" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 20:34Z 57 21 29 256 (WSW) 1055 - pcl 21:06Z 55 19 18 225 (SW) -9999 - " 21:38Z 59 20 18 252 (WSW) 1144 - " 01:32Z 52 20 17 267 (W) 2130 - " 02:04Z 51 18 21 246 (WSW) 1765 - " 02:37Z 50 17 21 257 (WSW) 1749 - " 03:08Z 49 17 17 257 (WSW) 1666 - " 03:39Z 50 17 20 235 (SW) 1810 - " 04:10Z 49 17 17 261 (W) 1669 - " 04:41Z 49 17 18 263 (W) 1494 - " 05:13Z 48 17 23 267 (W) 1398 - " 05:45Z 48 17 19 264 (W) 1394 - clear 06:17Z 47 17 24 263 (W) 1443 - pcl 06:50Z 46 17 22 268 (W) 1438 - clear 07:21Z 45 17 25 267 (W) 1459 - " 07:51Z 45 16 24 268 (W) 1458 - " 08:22Z 44 16 18 266 (W) 1438 - " 08:52Z 44 15 16 272 (W) 1552 - " 09:23Z 43 11 22 291 (WNW) 1453 - " 09:53Z 42 13 26 279 (W) 1377 - " 10:23Z 42 14 22 264 (W) 1427 - " 10:53Z 42 14 24 269 (W) 1428 - pcl 11:23Z 41 14 29 275 (W) 1431 - " pcl = partly cloudy ---=== TELESCOPE STATUS ===--- 02:40Z Enclosure doors opened, fans and blowers on, louvers opened. 02:45Z Enclosure off. 11:20Z Enclosure on, doors closed, fans and blowers off, louvers closed. Telescope stowed at 30 degrees with cartridge 2 (plate 1302) mounted. Counterweights at 280. Spectro auto-fill connected. Imager dewar 11:53Z 90 lb Spectro dewar 11:53Z 289 lb No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_113_1 Watcher: v2_22_0 MCP: v5_18_0 TPM: tpm_v2_28_0 AstroDa: v14_47 TCC: 2.6.8 sdssProcedures: v1_68 tccMon: v2_0 SoS: v4_9_13 hoggPT: v1_6_8 plate-mapper: v4_2_0 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0960 0.0790 Axial B 0.7970 0.7970 Axial C 0.8240 0.8240 Trans D -9.0670 -9.1130 Lateral E 1.8360 1.8870 Lateral F 0.0000 1.4300 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES PriDesOrient: 0.00 -11.80 23.00 1256.90 642.10 PriOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 13.9 Alt.: 29.999608 MIGS TONIGHT NOMINAL Axial A 1.5180 1.5170 Axial B 1.1550 1.1560 Axial C 1.0910 1.0990 Trans D -0.6870 -0.7070 GALILS Commanded: 1619649. 1555240. 1576354. -3400. -7900. Actual: 1619650. 1555250. 1576350. -3400. -7900. SETMIR VALUES SecDesOrient: 1257.00 0.00 -20.00 0.00 130.98 SecOrient: 1257.00 -0.02 -20.01 -0.48 130.89 ---=== PROBLEMS IN DETAIL ===---